Project

General

Profile

News

MICE News 13th May

Added by Kyberd, Paul over 5 years ago

News from the MOM

The first scheduled Machine Physics beam period for MICE on Tuesday evening was cancelled as ISIS
was still bringing up the machine.

- MICE and ISIS suffered a complete power cut last week on Tuesday. We lost power to the MICE Hall,
the rack room, and the MLCR at about noon. The reason for this outage is not yet known but is under
investigation. The DS was powered at 200A - it lost all He and began to warm up when the power cut out.
We refilled the He buffer, and the DS RAL team brought the magnet back to operational and it was turned
on again Friday morning. The single station tracker cryostat and electronics were also ON. The cryostat
lost the vac pumps but the automatic valve kicked in as designed when we lost power so recovery was
complete within the day. All racks, computers, servers, and IOCs in the Hall and target vac rack area
were recovered, rebooted, and regained full functionality by 7pm Tuesday. This includes access to the DS
mimics to aid in the magnet recovery.

- MICE took beam on Friday evening during our second scheduled shift. ISIS was up and running, and MICE
was ready. We successfully ran in the target with the adjusted lower limit on the frame and the
now-standard beam bump in ISIS. We are operating with a 2V loss limit. In order to increase this, we
will need to schedule an activation run during a Machine Physics period using the beam bump as all previous
studies were done with ISIS in the non-beam-bump configuration. Even with the limit, when we ran with a
272 MeV/c pion Reference Run beam, we were getting significant numbers of particle triggers with ~1000 in
TOF0 and ~450 in TOF1 when operating under the 2V limit. During running, we were also able to see triggers
in the single station tracker system at the same rate as in TOF1 (which was generating the trigger).

- The Tracker Controls rack was installed in the Hall last week with a prototype controls system that
will be used during the single station tracker test. The Tracker team continues to assemble at RAL with the
arrival of the Fermilab team on Saturday. Work continues on the tracker DAQ with a focus on timing in all
of the signals with the live time of the tracker electronics.

- A new Shifter Training regime has been instituted in the last week - contact the MOM for details if you
are scheduled to take any shifts in the near future. Even if you have previously been a shifter, you must
have your training assessed and be brought into the new system.

- MICE will be taking beam this week Tuesday through Sunday for the tracker single station, Software tests,
data for global tracking analysis, and Run Control integration.

- Stewart Greenall (MICE Hall Manager) is clearing out unnecessary equipment/cabinets/shipping crates/etc.
from the MICE Hall. If you (CKOV group, TOF group, etc) have ANYTHING in the Hall that you want to stay there,
email Stewart ().

(And finally)

Associate Professor

Yagmur Torun is to be promoted to associate professor with tenure at IIT.
Dan Kaplan believes he the first person in the world to get tenure based on muon-cooling R&D.
If you know better let us know.

MICE news 29th April

Added by Kyberd, Paul over 5 years ago

LH2 System

A little bit of news on the LH2 system progress. The vacuum problems have been solved and we
started the cooldown on Tuesday as scheduled. The system reached 4.2K last night and is now
beginning to condense helium for the first time in its final configuration.

Tracker single station test

- The single station tracker was fully installed in the Hall on Thursday by Geoff Barber, David Adey,
Ray Gamet, Ed Overton, Craig Macwaters, and Ken Long. The detector is mounted on the downstream end of
Q9, and a single tracker cryostat is located nearby. All lightguides for the single station test were
installed and now run from the detector to the cassettes on the cryostat. The cryostat is cold and the
DAQ is functioning well. Efforts to fully integrate the DAQ into the MLCR DAQ are ongoing with a focus
on timing in the light arrival time, the live time, and the TOF particle triggers.

- The Decay Solenoid is now cold and ready for powering. However, there are still some issues with the
chiller that will be worked on tomorrow. The DS will hopefully be in a good state for the first day of
data-taking on Tuesday 8th May.

- Dates of running during Machine Physics are now final - we have two shifts on 8 and 11 May. This will
be followed by running during the User Run on 15 May - 20 May. Requests for beam should be sent to the MOM.

- In preparation for running - a full test of all beamline magnets (except for the DS) and target was
performed last Wednesday and all looks good.

- Modeling of the magnetic field from the superconducting cooling channel magnets in Step IV around
sensitive material continues. This is necessary in order to develop specific plans to mitigate anticipated
problems with the equipment. Solving this problem for the cryocoolers along the south wall is currently
the priority. Anyone concerned about the fields near their equipment should contact Mike Courthold and
Vicky Bayliss.

- Preparation of R9 for the FC testing continues with final electrical work anticipated to begin soon
and installation of the chiller imminent.

- News on Spectrometer Solenoid progress and Analysis Group work will be given at the Videoconference
on Thursday - agenda soon.

News 15th April

Added by Kyberd, Paul almost 6 years ago

Our first collaboration paper Accepted

The beamline paper submitted to JINST has been accepted with minor modifications.

Field Mapping Equipment

Pierre Ange informs us (message below) that the mapping equipment was shipped from CERN last Friday! All we need now is magnets!

"_Par ce présent courrier, je t'informe que la caisse(mapping Equipment) est parti Aujourd'hui.
Elle devrait arrivée courant de la semaine prochaine. J'ai informé Andy.
Par ailleurs, aurais-tu les dates un peu plus précises pour nous?
En espérant te trouver en forme, je te souhaite un bon W.E. et a bientôt !
Cordiales salutations.
Pierre-ange GIUDICI_"

News Addendum: March Running

Added by Kyberd, Paul almost 6 years ago

MICE March Run

MICE ran from the evening of Thursday 22nd March through to Sunday 25th March.
Unfortunately ISIS had a major power outage during Sunday, preventing further operation,
although most of the aims of the run had already been achieved.

Most of the magnets and detectors in the beamline functioned well with no major issues to report.
The Decay Solenoid was not operated due to an imminent power supply upgrade. The DAQ was run on MICERaid1
as MICERaid 2 requires repair.

A set of target scans were undertaken and completed. These were done to see if there was any notable
difference in the beam position/beam bump since the last run. Further analysis will be required to
see if these scans show any difference

A scan of D1 was completed, for a 200Mev/ pion/muon beam. The data awaits further analysis.
A measurement of the beam timing with respect to the ISIS RF clock was completed

Tests were made on the reconstruction software. Good progress was made although with ISIS down
some final runs to check the reconstruction and the CKOVs was not possible, although
of the Cerenkov online monitoring was demonstrated.

Some work was done on the improved run controls but due to the power cut not all the tests
were completed.

Week ending 18th March

Added by Kyberd, Paul almost 6 years ago

Message from Alain Blondel

The MICE project board took place on 8 March. The slides that were shown are here
https://indico.cern.ch/conferenceDisplay.py?confId=180596.
The preliminary close-out report from the board (quite positive) can be found there too.
The MICE report to MPB was produced well on time (we are improving!) and can be found as
MICE note 366 (edited in great part by Dan Kaplan).
Thanks to all MICE for this effort!

Message from the EB

Having shipments planned well in advance helps avoid unexpected and unintentional costs.
Please notify Andy Nichols well in advance if you are shipping something to RAL.

Our first collaboration paper

"The MICE Muon Beam on ISIS and the beam-line instrumentation of the Muon Ionization Cooling Experiment"
has been finally submitted to arXiv (temporary submission identifier:
submit/0438465, scheduled announcement at Tue, 20 Mar 2012 00:00:00 GMT),
thanks to the efforts of many mouses, including Paul, Ken, Dan, Alain, Marco ...
Next week it will be submitted also to JINST.

Building work in the MICE Hall

Added by Kyberd, Paul almost 6 years ago

Building work in the MICE hall

Please note that there will be building work (fireproofing) around the MICE Hall personnel entrance
and metal detector, starting on Monday 30 Jan. for up to two weeks.
During this period there will be no access into the Hall via that route.
Hats and vests have been moved round to the equipment door.

Air conditioning work in the MICE hall

Scaffolding has already been installed in the Hall for the air-conditioning reconfiguration
- units 2,3 and 4 will have their baffles and fans modified, and unit 5 will be moved on to
the east end of the LH2 mezzanine.
This process is expected to start on Feb 6th and be complete by the 29th February; in the
meantime various magnet and HV supplies may need to be turned off.

First MICE news of the year

Added by Kyberd, Paul about 6 years ago

A Christmas Mouse

With thanks to Gail Hanson and apologies for late posting (I have had some network problems).

Limited Parking at Ridgeway

Please be advised that the construction of 63 new car parking spaces, to replace the existing 10 spaces
at R87 Ridgeway House will commence as of the 03.01.12. Work is expected to take approximately 8 weeks,
although at this time of year the weather may affect this programme, especially given that he whole project
consists of ground works. Some flexibility in project completion dates may therefore be required.
The work is to be undertaken in 2 phases, with the first concentrating on the area between R31 and R87 Wing 2.  
Once this is complete, the road and existing 10 car parking spaces will be remodelled to create the final 22
spaces, including 2 for disabled people.

MICE News 12 December 2011

Added by Kyberd, Paul about 6 years ago

Data Volume

As of run 3426, MICE has now taken more than a terabyte of data during Step 1!

MOM update

So far we have very successfully data taking period.<p>

Preliminary calibration of the TOF0 and TOF1 detectors shows that after
the refurbishment of TOF1 we have a significant improvement in the
resolution this station. Now both TOF0 and TOF1 have resolution of ~50 ps.

Progress has been made toward the understanding of the electron/positron
different time-of-flight puzzle. It has been demonstrated that variation
of the number of particle trigger per spill introduces difference in the
measured time-of-flight even when we use identical settings of the
beamline channel.

Limited Restaurant Service at RAL

On Tuesday 13 and Wednesday 14 December the Restaurant will be open for pre-booked Christmas lunches only.
The R22 coffee lounge will be open as normal. On 13 and 14 December the evening meal service will run as normal.

Alternative services:
A brunch service will be available from R22 from 07:30 to 11:00 (on 13 and 14 December)
The R22 Deli bar will be open as normal serving baguettes, rolls, sandwiches, Panini, soup,
jacket potatoes and a take away pasta dish (on 8, 13 and 14 December)
The R1 Coffee Lounge will be open as normal from 08:30 to 15:30 and provides a range of snacks.
Soup and jacket potatoes will also be available.
The Electron Cafe will be open from 08:30 to 15:30 and provides a range of snacks, one hot meal
during the lunch-time and Costa Coffee.

(81-90/114)

Also available in: Atom